# sPhinX: Sample Efficient Multilingual Instruction Fine-Tuning Through N-shot Guided Prompting

### v2

@@ A.2 Detailed Results
Language en fr jp pt ru zh avg
Mistral-7B
Base Model 0.52 0.47 0.52 0.54 0.54 0.50 0.52
IFT 0.61 0.57 0.57 0.57 0.60 0.56 0.58
m-Alpaca 0.61 0.57 0.57 0.57 0.60 0.56 0.58
Aya 0.55 0.56 0.54 0.54 0.56 0.54 0.55
Bactrian 0.61 0.57 0.57 0.57 0.60 0.56 0.58
sPhinX-0s 0.75 0.65 0.68 0.67 0.66 0.65 0.68
sPhinX 0.80 0.69 0.72 0.70 0.67 0.67 0.71
Phi-3-small
Base Model 0.86 0.67 0.73 0.77 0.74 0.72 0.75
IFT 0.86 0.78 0.72 0.78 0.77 0.75 0.78
m-Alpaca 0.87 0.76 0.75 0.78 0.76 0.71 0.81
Aya 0.79 0.61 0.67 0.70 0.70 0.66 0.69
Bactrian 0.83 0.72 0.71 0.75 0.70 0.68 0.73
sPhinX-0s 0.89 0.77 0.78 missing0.79 0.81 0.76 0.80
sPhinX missing0.89 missing0.76 missing0.79 0.78 missing0.82 missing0.77 missing0.84
Language ar de el en es hi ro ru th tr vi zh avg
Mistral-7B
Base Model 0.62 0.81 0.64 0.89 0.86 0.65 0.82 0.71 0.59 0.68 0.79 0.76 0.74
IFT 0.42 0.68 0.33 0.92 0.66 0.5 0.71 0.61 0.38 0.63 0.71 0.68 0.60
Aya 0.33 0.73 0.65 0.85 0.8 0.63 0.75 0.67 0.57 0.61 0.75 0.59 0.66
sPhinX-0s 0.54 0.76 0.7 0.88 0.84 0.69 0.77 0.66 0.52 0.64 0.71 0.60 0.69
sPhinX 0.74 0.87 0.76 0.93 0.90 0.79 0.86 0.77 0.63 0.76 0.88 0.73 0.80
Phi-3-small
IFT 0.71 0.88 0.73 0.92 0.91 0.64 0.84 0.80 0.44 0.70 0.67 0.76 0.75
Aya 0.61 0.89 0.84 0.94 0.93 0.80 0.89 0.82 0.73 0.83 0.92 0.79 0.83
Bactrian 0.81 0.92 0.81 0.95 0.95 0.80 0.91 0.84 0.74 0.82 0.92 0.79 0.85
Language et ht id it qu sw ta th tr vi zh en avg
Mistral-7B
IFT 0.52 0.52 0.69 0.79 0.50 0.51 0.50 0.54 0.57 0.63 0.75 0.90 0.62
Aya 0.57 0.54 0.64 0.67 0.53 0.56 0.57 0.62 0.56 0.61 0.64 0.78 0.61
sPhinX-0s 0.54 0.5 0.58 0.63 0.51 0.55 0.52 0.52 0.54 0.57 0.64 0.8 0.58
sPhinX missing0.64 0.54 0.73 0.80 0.53 0.61 0.59 0.63 0.67 0.66 0.80 0.91 0.68
Phi-3-small
IFT 0.55 0.57 0.81 0.93 0.53 0.58 0.48 0.60 0.62 0.69 0.88 0.96 0.68
Aya 0.60 0.55 0.72 0.83 0.52 0.55 0.52 0.62 0.59 0.69 0.75 0.89 0.65
Bactrian 0.62 0.56 0.83 0.91 0.54 0.62 0.52 0.66 0.65 0.71 0.86 0.99 0.71
sPhinX-0s 0.59 0.58 0.85 0.94 0.50 0.60 0.54 0.63 0.69 missing0.72 0.89 0.96 0.71
sPhinX 0.59 missing0.60 missing0.85 missing0.94 missing0.52 missing0.57 missing0.58 missing0.68 missing0.68 0.71 missing0.90 missing0.98 missing0.72
Language ar en es eu hi id my ru sw te zh avg
Mistral-7B
Base Model 0.65 0.89 0.83 0.56 0.62 0.76 0.52 0.81 0.56 0.52 0.80 0.68
IFT 0.70 0.96 0.92 0.54 0.69 0.79 0.57 0.90 0.58 0.54 0.88 0.73
m-Alpaca 0.53 0.73 0.7 0.51 0.51 0.57 0.50 0.66 0.52 0.52 0.71 0.59
Aya 0.64 0.86 0.81 0.56 0.71 0.73 0.60 0.82 0.67 0.60 0.81 0.71
Bactrian 0.69 0.82 0.74 0.52 0.59 0.76 0.54 0.73 0.62 0.61 0.76 0.67
sPhinX-0s 0.57 0.66 0.64 0.47 0.56 0.61 0.50 0.62 0.56 0.52 0.69 0.58
sPhinX 0.83 0.96 0.94 0.57 missing0.84 0.87 missing0.61 0.91 missing0.80 missing0.69 0.94 0.81
Phi-3-small
Base Model 0.80 0.98 0.96 0.61 0.72 0.92 0.53 0.96 0.61 0.55 0.94 0.78
IFT 0.81 0.98 0.96 0.61 0.75 0.92 0.56 0.96 0.61 0.53 0.94 0.79
m-Alpaca 0.81 0.99 0.98 0.58 0.76 0.93 0.52 0.97 0.64 0.54 0.96 0.79
Aya 0.77 0.98 0.97 0.57 0.77 0.93 0.53 0.96 0.74 0.56 0.94 0.79
Bactrian 0.83 0.98 0.98 0.61 0.83 0.94 0.54 0.97 0.79 0.63 0.94 0.82
Language ar de es en fi fr hi it jp ko ta te vi zh avg
Mistral-7B
Phi-3-small
Model XC XS XW XQ BL
Mistral-7B
Base Model 0.63 0.68 0.52 0.74 0.24
IFT 0.62 0.73 0.54 0.60 0.47
m-Alpaca 0.55 0.59 0.51 0.46 0.41
Aya 0.68 0.71 0.54 0.66 0.38
Bactrian 0.54 0.67 0.54 0.69 0.26
sPhinX-0s 0.58 0.58 0.68 0.69 0.67
sPhinX 0.68 0.81 0.71 0.80 0.71
Phi-3-small
Base Model 0.64 0.78 0.75 0.78 0.65
IFT 0.68 0.79 0.78 0.75 0.70
m-Alpaca 0.68 0.79 0.81 0.77 0.75
Aya 0.65 0.79 0.69 0.83 0.72
Bactrian 0.71 0.82 0.73 0.85 0.77
sPhinX-0s 0.71 0.81 0.80 0.82 0.79
sPhinX 0.72 0.84 0.87 0.87 0.79
Table 9: Performance of Mistral-7B and Phi-3-small instruction-tuned for 10000 training steps on various datasets. Abbreviations: XC - XCOPA, XS - XStoryCloze, XW - XWinograd, XQ - XQuAD, BL - Belebele. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Benchmarks Base Model sPhinX
MMLU
(5-shot)
0.76 0.75
HellaSwag
(5-shot)
0.81 0.83
GSM-8k
(8-shot, CoT)
0.85 0.77
MedQA
(2-shot)
0.64 0.66
Arc-C
(10-shot)
0.90 0.90
Arc-E
(10-shot)
0.97 0.97
PIQA
(5-shot)
0.84 0.89
WinoGrande
(5-shot)
0.77 0.82
OpenBookQA
(10-shot)
0.86 0.88
BoolQ
(2-shot)
0.82 0.87
CommonSenseQA
(10-shot)
0.80 0.81

### v4

@@ A.5 Detailed Results
[…]
Translate the following sentence pairs:
{Source Language}: {Source Phrase} {Target Language}: {Target Phrase}
Figure 10: Translation evaluation prompt
[…]
System: You are a helpful following assistant whose goal is to select the preferred (least wrong) output for a given instruction in {LANGUAGE_NAME}.
User: Which of the following answers is the best one for given instruction in {LANGUAGE_NAME}. A good answer should follow these rules: 1) It should be in {[LANGUAGE_NAME}.
2) It should answer the request in the instruction.
3) It should be factually and semantically comprehensible.
4) It should be grammatically correct and fluent.
Instruction: {INSTRUCTION}
Answer (A): {COMPLETION A}
Answer (B): {COMPLETION B}
FIRST provide a one-sentence comparison of the two answers, explaining which you prefer and why. SECOND, on a new line, state only ‘Answer (A)’ or ‘Answer (B)’ to indicate your choice. If the both answers are equally good or bad, state ‘TIE’. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
Preferred: <‘Answer (A)’ or ‘Answer (B)’ or ‘TIE’>
Figure 11: Preference simulation prompt taken from (Üstün et al., 2024) evaluation suite to evaluate our models on free-form generation using GPT-4o.
@@
Dataset Average Token
Length/Sample
Aya 2240
Bactrian 2465
m-Alpaca 1620
sPhinX-0s 544
sPhinX 3100
Table 6: Average Token Length in each dataset
@@
Figure 12: Some examples of input queries and its counterpart existing in the hindi version of the Multialpaca dataset and if it was generated using the Selective Translated Augmentation strategy. Again, we observe that the samples generated using Selective Translated Augmentation translate only the required amount of information as controlled via prompting whereas in Multialpaca the translations are direct translations where only a part of the instructions have been followed to translate the input queries.
Language en fr jp pt ru zh avg
Mistral-7B
Base Model 0.52 0.47 0.52 0.54 0.54 0.50 0.52
IFT 0.61 0.57 0.57 0.57 0.60 0.56 0.58
m-Alpaca 0.61 0.57 0.57 0.57 0.60 0.56 0.58
Aya 0.55 0.56 0.54 0.54 0.56 0.54 0.55
Bactrian 0.61 0.57 0.57 0.57 0.60 0.56 0.58
sPhinX-T 0.58 0.61 0.57 0.53 0.54 0.57 0.57
sPhinX-0s 0.75 0.65 0.68 0.67 0.66 0.65 0.68
sPhinX 0.80 0.69 0.72 0.70 0.67 0.67 0.71
Phi-3-small
Base Model 0.86 0.67 0.73 0.77 0.74 0.72 0.75
IFT 0.86 0.78 0.72 0.78 0.77 0.75 0.78
m-Alpaca 0.87 0.76 0.75 0.78 0.76 0.71 0.81
Aya 0.79 0.61 0.67 0.70 0.70 0.66 0.69
Bactrian 0.83 0.72 0.71 0.75 0.70 0.68 0.73
sPhinX-T 0.87 0.74 0.74 0.77 0.77 0.72 0.77
sPhinX-0s 0.88 0.75 0.78 missing0.79 0.81 0.76 0.80
sPhinX missing0.89 missing0.76 missing0.79 missing0.79 missing0.82 missing0.77 missing0.84
Language ar de el en es hi ro ru th tr vi zh avg
Mistral-7B
Base Model 0.62 0.81 0.64 0.89 0.86 0.65 0.82 0.71 0.59 0.68 0.79 0.72 0.73
IFT 0.42 0.68 0.33 0.92 0.66 0.5 0.71 0.61 0.38 0.63 0.71 0.68 0.60
Aya 0.33 0.73 0.65 0.85 0.80 0.63 0.75 0.67 0.57 0.61 0.75 0.59 0.66
sPhinX-T 0.71 0.83 0.75 0.92 0.89 0.77 0.84 0.75 0.60 0.77 0.86 0.63 0.78
sPhinX-0s 0.54 0.76 0.70 0.88 0.84 0.69 0.77 0.66 0.52 0.64 0.71 0.60 0.69
sPhinX 0.74 0.87 0.77 0.93 0.90 0.79 0.86 0.77 0.63 0.77 0.88 0.73 0.80
Phi-3-small
IFT 0.71 0.88 0.73 0.92 0.91 0.64 0.84 0.80 0.44 0.70 0.67 0.76 0.75
Aya 0.61 0.89 0.84 0.94 0.93 0.80 0.89 0.82 0.73 0.83 0.91 0.79 0.83
Bactrian 0.81 0.92 0.81 0.95 0.95 0.80 0.90 0.84 0.72 0.82 0.91 0.79 0.85
sPhinX-T 0.80 0.91 0.82 0.95 0.94 0.80 0.90 0.83 0.69 0.81 0.91 0.73 0.78
Language et ht id it qu sw ta th tr vi zh en avg
Mistral-7B
IFT 0.52 0.52 0.69 0.79 0.50 0.51 0.50 0.54 0.57 0.63 0.75 0.90 0.62
Aya 0.57 0.54 0.64 0.67 0.53 0.56 0.57 0.62 0.56 0.61 0.64 0.78 0.61
sPhinX-T 0.57 0.50 0.64 0.72 0.50 0.50 0.58 0.57 0.57 0.62 0.71 0.83 0.61
sPhinX-0s 0.54 0.5 0.58 0.63 0.51 0.55 0.52 0.52 0.54 0.57 0.64 0.8 0.58
sPhinX missing0.64 0.54 0.73 0.80 0.53 missing0.61 0.59 0.63 0.67 0.66 0.80 0.91 0.68
Phi-3-small
IFT 0.55 0.57 0.81 0.93 0.53 0.58 0.48 0.60 0.62 0.69 0.88 0.96 0.68
Aya 0.60 0.55 0.72 0.83 0.52 0.55 0.52 0.62 0.59 0.69 0.75 0.89 0.65
Bactrian 0.62 0.56 0.83 0.91 0.52 0.60 0.52 0.66 0.65 0.71 0.86 0.98 0.70
sPhinX-T 0.55 0.58 0.83 0.93 0.51 0.57 0.56 0.66 0.68 0.68 0.87 0.98 0.70
sPhinX-0s 0.59 0.58 0.84 0.93 0.50 0.60 0.54 0.63 0.68 missing0.72 0.89 0.96 0.71
sPhinX 0.59 missing0.60 missing0.85 missing0.94 missing0.52 0.57 missing0.58 missing0.68 missing0.69 0.71 missing0.90 missing0.99 missing0.72
Language ar en es eu hi id my ru sw te zh avg
Mistral-7B
Base Model 0.65 0.89 0.83 0.56 0.62 0.76 0.52 0.81 0.56 0.52 0.80 0.68
IFT 0.70 0.95 0.92 0.54 0.69 0.79 0.57 0.90 0.58 0.54 0.88 0.73
m-Alpaca 0.53 0.73 0.70 0.51 0.51 0.57 0.50 0.66 0.52 0.52 0.71 0.59
Aya 0.64 0.86 0.81 0.56 0.71 0.73 0.60 0.82 0.67 0.60 0.81 0.71
Bactrian 0.69 0.82 0.74 0.52 0.59 0.76 0.54 0.73 0.62 0.61 0.76 0.67
sPhinX-T 0.78 0.92 0.87 0.56 0.80 0.81 0.66 0.86 0.74 0.70 0.86 0.78
sPhinX-0s 0.57 0.66 0.64 0.47 0.56 0.61 0.50 0.62 0.56 0.52 0.69 0.58
sPhinX 0.83 0.96 0.94 0.57 missing0.84 0.87 missing0.67 0.91 missing0.80 missing0.69 0.94 0.81
Phi-3-small
Base Model 0.80 0.98 0.96 0.61 0.72 0.92 0.53 0.96 0.61 0.55 0.94 0.78
IFT 0.81 0.98 0.96 0.61 0.75 0.92 0.56 0.96 0.61 0.53 0.94 0.79
m-Alpaca 0.81 0.98 0.98 0.58 0.76 0.93 0.52 0.97 0.64 0.54 0.96 0.79
Aya 0.77 0.98 0.97 0.57 0.77 0.93 0.53 0.96 0.74 0.56 0.94 0.79
Bactrian 0.83 0.98 0.98 0.61 0.83 0.94 0.54 0.97 0.79 0.63 0.94 0.82
sPhinX-T 0.84 0.98 0.98 0.60 0.80 0.95 0.52 0.96 0.72 0.60 0.85 0.80
Language ar de es en fi fr hi it jp ko ta te vi zh avg
Mistral-7B
sPhinX-T 0.66 0.74 0.71 0.81 0.66 0.76 0.57 0.73 0.66 0.68 0.54 0.46 0.66 0.71 0.68
Phi-3-small
sPhinX-T 0.71 0.90 0.90 0.93 0.73 0.90 0.56 0.89 0.82 0.75 0.42 0.38 0.75 0.87 0.75
Language ar fr de ro ja ru zh avg
Mistral-7B
Base Model 0.48 0.63 0.65 0.56 0.43 0.54 0.48 0.54
IFT 0.37 0.61 0.61 0.58 0.28 0.5 0.49 0.49
m-Alpaca 0.34 0.51 0.51 0.46 0.29 0.36 0.41 0.41
Aya 0.30 0.50 0.51 0.46 0.24 0.35 0.41 0.39
Bactrian 0.40 0.55 0.52 0.51 0.34 0.44 0.42 0.45
sPhinX-T 0.39 0.60 0.60 0.54 0.39 0.49 0.48 0.49
sPhinX-0s 0.40 0.57 0.61 0.53 0.40 0.51 0.44 0.49
sPhinX 0.45 0.63 0.64 0.59 0.45 0.55 0.52 0.54
Phi-3-small
Base Model 0.48 0.61 0.65 0.57 0.45 0.52 0.53 0.54
IFT 0.43 0.62 0.63 0.52 0.41 0.49 0.50 0.54
m-Alpaca 0.44 0.60 0.61 0.56 0.32 0.44 0.19 0.45
Aya 0.44 0.56 0.57 0.54 0.12 0.45 0.17 0.41
Bactrian 0.48 0.63 0.65 0.59 0.45 0.52 0.52 0.54
sPhinX-T 0.48 0.64 0.65 0.59 0.46 0.53 0.52 0.55
sPhinX-0s 0.49 0.63 0.65 0.59 0.46 0.54 0.53 0.56
sPhinX 0.49 0.64 0.66 0.60 0.46 0.54 0.53 0.56
Table 13: Granular results for Translation for language to English direction (4-shot) on our model. Metric: ChrF. The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
Language ar fr de ro ja ru zh avg
Mistral-7B
Base Model 0.29 0.60 0.54 0.52 0.21 0.34 0.47 0.42
IFT 0.16 0.58 0.57 0.48 0.17 0.29 0.43 0.38
m-Alpaca 0.15 0.62 0.66 0.40 0.12 0.48 0.44 0.41
Aya 0.12 0.54 0.63 0.48 0.16 0.39 0.42 0.39
Bactrian 0.14 0.51 0.49 0.44 0.10 0.35 0.40 0.38
sPhinX-T 0.31 0.58 0.53 0.47 0.20 0.30 0.26 0.38
sPhinX-0s 0.30 0.60 0.55 0.51 0.22 0.31 0.45 0.42
sPhinX 0.35 0.61 0.60 0.55 0.26 0.36 0.49 0.46
Phi-3-small
Base Model 0.31 0.63 0.60 0.43 0.24 0.30 0.45 0.42
IFT 0.29 0.61 0.58 0.39 0.21 0.27 0.41 0.46
m-Alpaca 0.39 0.60 0.58 0.31 0.11 0.24 0.47 0.38
Aya 0.17 0.62 0.56 0.45 0.22 0.28 0.46 0.39
Bactrian 0.30 0.60 0.56 0.47 0.18 0.24 0.44 0.40
sPhinX-T 0.33 0.63 0.60 0.48 0.24 0.31 0.48 0.43
sPhinX-0s 0.32 0.63 0.61 0.50 0.27 0.36 0.49 0.45
sPhinX 0.35 0.64 0.61 0.51 0.26 0.33 0.49 0.46
Table 14: Granular results for Translation for English to language direction (4-shot) on our model. Metric: ChrF. The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.